OpenStudy (anonymous):

If x*x + 2x - 35 = 0, then x = ? please help

OpenStudy (zpupster):

factor the polynomial (x )(x ) can you do this step??

OpenStudy (anonymous):

NOPE

OpenStudy (zpupster):

ok (x )(x ) we need to find factors of 35, we will deal with the signs later so factors are 1,5,7,35 the we look at what factors(the difference will = 2( the middle term) 7 and 5 look like a good choice because the difference between those numbers is 2 (x 7 )(x 5 ) now we have to find the signs since we have -35 we know that a pos and neg = a neg and we have +2 so the bigger +7-5=+2 (x +7 )(x-5 )=0 now we set both of these binomial to zero x+7=0 x-5=0 x+7-7=0-7 we subtract 7 both sides one x=-7 x-5=0 here you can do the same by adding 5 to both sides to get the other answer

OpenStudy (zpupster):

when you get the 2 answers plug it back in the original and see if they are correct